# wd1: Weighted Policy Optimization for Reasoning in Diffusion Language Models

### v1

@@ 3.2 Method
4:     Sample prompt $q\sim\mathcal{D}$
5:     Sample $G$ completions $o_{i}\sim\pi^{\text{ref}}_{{\text{old}}}(\cdot\mid q),\forall i\in[G]$
6:     Compute group-relative advantage $\hat{A}_{i}=R(q,o_{i})-\texttt{mean}(R(q,o_{1:G}))$, $\forall i\in[G]$
7:     Compute weights $w^{+}$ and $w^{-}$ in Equation 9, $\forall i\in[G]$
8:     for gradient update iterations $n=1,2,\dots,\mu$  do
9:        Compute approximation of log-likelihood $\log\pi_{\theta}(o_{i}|q)$
10:        Compute objective $\mathcal{L}_{\texttt{wd1}}(\theta)$ in Equation 8
11:        Update $\pi_{\theta}$ by gradient descent
12:     end for
13:  end while
14:  return  $\pi_{\theta}$
The objective in Equation 7 aims to amplify the likelihood of completions with high advantage. Crucially, due to the exponential form of the weighting in $\mathcal{L}_{NLL}$, completions with relatively low advantage – referred to as negative samples – may receive vanishingly small weights. To address this imbalance and make full use of the samples, we propose a complementary penalty term that minimizes the likelihood of low-advantage completions. Building on the weighted negative log-likelihood in Equation 7, we propose a novel weighted log-likelihood objective for dLLM policy optimization, where the weights are determined by group-relative advantage. Specifically, we define the original weights as $w^{+}$ and introduce negative weights $w^{-}$ to downweight the likelihood of completions with low advantage:
where the weights are defined with group-relative (GRPO) advantage $\hat{A}_{i}=R(q,o_{i})-\texttt{mean}(R(q,o_{1:G}))$:
Our method wd1, the first weighted policy optimization for dLLMs, is formally presented in Algorithm 1. We first obtain $G$ completions $\{o\}_{i=1}^{G}$ sampled from geometric mixture $\pi_{\text{old}}^{\text{ref}}(\cdot|q)\propto\pi_{\text{old}}(\cdot|q)^{\lambda/(\lambda+\beta)}\cdot\pi_{{\text{ref}}}(\cdot|q)^{\beta/(\lambda+\beta)}$ (line 5). Since our base model LLaDA parametrizes the clean token prediction $\pi^{\text{ref}}_{\text{old}}(x_{0}^{k}|x_{t},q)$ for denoising, we approximate $\log\pi^{\text{ref}}_{\text{old}}(x_{0}^{k}|x_{t},q)\approx\lambda\log\pi_{\text{old}}(x_{0}^{k}|x_{t},q)+\beta\log\pi_{\text{ref}}(x_{0}^{k}|x_{t},q)$ as the logits of the denoising distribution at each step $t$. We then use the samples to compute weights in Equation 9 (line 6). In weights computing, we leverage estimated partition functions $\sum_{j=1}^{G}\exp\big{(}A(q,o_{j})\big{)}$ and $\sum_{j=1}^{G}\exp\big{(}-A(q,o_{j})\big{)}$ for normalization, in order to restrict the absolute value of loss and thus restricting the gradient norm and stabilizing the training. Finally, we leverage the likelihood approximation introduced in d1: $\log\pi_{\theta}(o_{i}|q)\approx\sum_{k}\log\pi_{\theta}(o_{i}^{k}|q^{\prime})$ (line 8), where $q^{\prime}$ is randomly masked from $q$ at every gradient step $n$.
We simplify policy optimization to a weighted likelihood formulation, such that, during reasoning with dLLMs, the only required likelihood approximation is that of the current policy $\pi_{\theta}$. This design significantly improves computational efficiency while reducing training bias and instability. In contrast to prior methods that rely extensively on multiple likelihood approximations across several policies, wd1 offers a more streamlined and scalable reinforcement learning approach. As a result, wd1 serves as a simple yet effective framework for scaling reasoning with diffusion-based language models [^1] In wd1 we include an additional term dependent on $w^{-}$ in our objective. This make our objective different from Equation 6 and may alter the monotonic guarantee. We leave the theoretical analysis in the future work..
@@ 4 Experiments
    Improved reasoning capabilities and higher accuracy than existing method on popular reasoning benchmarks without SFT;
@@
    reduced computational burden, as reflected by decreased runtime, lower FLOPs and numbers of function evaluations (NFEs) of the model; and
@@
Experimental Setup. We perform reinforcement learning (RL) fine-tuning on the LLaDA-8B-Instruct model (Nie et al., 2025a), on four different benchmarks: GSM8k (Cobbe et al., 2021), MATH (Lightman et al., 2023), Sudoku (Arel, 2025), and Countdown (Pan et al., 2025). Our main baseline is d1 (Zhao et al., 2025), the first RL method developed for masked diffusion LLMs (dLLMs). To ensure a fair comparison with our main baseline models, we use the same training dataset as in d1 (Zhao et al., 2025). [^2] We note that other concurrent works use larger datasets, which may lead to different performance outcomes on the same benchmarks (Yang et al., 2025; Zhu et al., 2025a; Huang et al., 2025). Specifically, for GSM8k and MATH, we use the respective training splits, while for Sudoku and Countdown, we adopt the dataset splits provided by Zhao et al. (2025). We use s1K  (Muennighoff et al., 2025) data for SFT in d1. We report results using zero-shot evaluation on sequence lengths of 256 and 512 tokens, with performance recorded at significantly less gradient steps than d1. Following d1, we report results from the best-performing checkpoint (more details provided in Appendix B.1).
Implementation. For a fair comparison, we reproduce the baseline methods Diffu-GRPO, which applies diffusion-based GRPO training directly to the LLaDA base model, and d1, which performs SFT before applying Diffu-GRPO. In our implementation of wd1, we apply the same likelihood approximation method as d1. Since previous works (Yu et al., 2025) have demonstrated that the reference policy is empirically unnecessary, we set $\beta=0$ and $\lambda=1$ to eliminate $\pi_{\text{ref}}$ in practice. The hyperparameters used in our method and our reproduction of d1 are listed in Table 5. The hyperparameters of SFT are provided in Table 4. For both methods, RL training is conducted on four NVIDIA A100 GPUs (80GB), and SFT is performed on four A6000 GPUs (48GB).
Model Sudoku Countdown GSM8K MATH500
@@
wd1 25.2 24.2 51.2 46.1 80.8 82.3 34.4 39.0
Table 1: Test Accuracy (%) across Different Tasks. wd1, our approach without SFT, demonstrates consistently higher accuracy, particularly significant on Sudoku and Countdown.
Model SFT RL Training
@@
wd1 0 hrs 82.16 sec/step $8.887\times 10^{15}$/step $\mu$/step
Table 2: Comparison of Training Cost on 4$\times$×A100. We show SFT cost, average training time, FLOPs evaluated by DeepSpeed Flops Profiler, and theoretical NFEs per global step which includes $\mu=8$ gradient steps. wd1 eliminates the need for SFT and has less cost per-global-step in RL compared to d1.
@@ 4.1 Main Results
Superior Reasoning Ability. In Table 1, we observe that wd1, even without supervised fine-tuning or using any supervised data, consistently outperforms our reproduced implementation of d1. Notably, wd1 surpasses d1 by 8% in test accuracy on the Sudoku task, and achieves up to a 25% improvement on Countdown with maximum length 256. Relative to the base LLaDA model, the performance gain reaches as high as 38%. On math problem-solving benchmarks such as GSM8K and MATH500, our approach attains slightly higher accuracy, further underscoring the effectiveness of wd1 across tasks.
Reduced Training Cost. Table 2 demonstrates that the training cost required by wd1 is substantially lower than that of d1. Unlike d1, wd1 does not require a SFT stage, which alone accounts for approximately two hours of training in d1. wd1 achieves additional speedup during the RL phase, where runtime is measured by averaging over $\mu=8$ inner gradient steps per global step. Notably, the time efficiency gap is expected to widen further under settings with larger maximum sequence lengths and more diffusion steps. This efficiency gain is further supported by a reduced FLOPs and number of function evaluations (NFEs) per step, as wd1 bypasses the need to approximate the likelihood of the old policy. We exclude NFEs associated with sampling, since both methods share identical sampling costs as wd1 removes the reference policy regularization.
4.2 Training Dynamics
Figure 2: Training Rewards Dynamics of wd1 and d1. Standard deviation reported over a rolling window of 50 steps. The average reward of the samples generated by wd1 increase faster than the baseline d1.
Figure 2 presents the reward dynamics over gradient steps during training. wd1 exhibits a notably faster reward increase compared to d1, highlighting its superior sample efficiency–effectively leveraging the reward signal to accelerate policy optimization. In addition, Figure 4 shows the average length of generated completions during training. On math reasoning benchmarks such as GSM8K and MATH500, wd1 converges to shorter output sequences than d1, suggesting improved token efficiency while maintaining or improving performance.
@@ 4.3 Ablation Study
Model Sudoku Countdown GSM8K MATH500
@@
wd1-SFT 26.5 24.2 43.4 43.4 80.7 82.0 36.4 39.0
wd1-P 6.69 6.84 13.67 4.69 65.66 78.17 29.40 22.80
Table 3: Ablation on SFT and Negative Samples Weight ($w^{-}$), denoted by wd1-SFT and wd1-P (positive only), respectively. Results show that wd1 performs well without SFT, supporting R1-Zero-style training. Removing negative sample reinforcement ($w^{-}$) significantly hurts performance, highlighting its importance.
We present an ablation study in Figure 3. Notably, we observe that supervised fine-tuning (SFT) yields only marginal improvements within our approach, with a slight gain in the Sudoku task. This contrasts with d1, where SFT plays a significant role in improving performance. These findings indicate that wd1 can eliminate the need for an SFT phase, thereby simplifying the training pipeline and substantially reducing computational cost. Additionally, we evaluate the impact of removing the negative-weighted term by setting $w^{-}=0$, thus relying solely on the positive advantage weights $w^{+}$. The resulting degradation in performance highlights the importance of explicitly penalizing the likelihood of completions with low advantage, reinforcing the contribution of negative samples in effective policy optimization.

### v2

@@ 3.1 Reinforcement Learning as Weighted Log-Likelihood Maximization
4:  Sample prompt $q\sim\mathcal{D}$ and $G$ completions $o_{i}\sim\pi_{{\text{old}}}(\cdot\mid q),\forall i\in[G]$
5:  Compute advantage $\hat{A}_{i}=R(q,o_{i})-\texttt{mean}(R(q,o_{1:G}))$, $\forall i\in[G]$
6:  Compute weights $w^{+}$ and $w^{-}$ in Equation˜9, $\forall i\in[G]$
7:  for gradient update iterations $n=1,2,\dots,\mu$ do
8:   Compute approximated log-likelihood $\log\pi_{\theta}(o_{i}|q)$
9:   Compute objective $\mathcal{L}_{\textit{wd1}}(\theta)$ in Equation˜8 or Equation˜10 and update $\theta$
10:  end for
11:end while
12:return $\pi_{\theta}$
4 Theoretical Insights: Energy-Guided Diffusion Sampling
In this section, we present a novel theoretical interpretation of policy optimization for dLLMs. We prove that the advantage-weighted log-likelihood objective (wd1) for dLLMs can be viewed as energy-guided discrete diffusion training combined with negative sample unlearning.
Sampling from the solution policy of the reverse-KL policy optimization, as described in Equation˜5, can be interpreted as energy-guided sampling, where the energy function $\mathcal{E}(q,\cdot)=-{A^{\pi_{\text{old}}}(q,\cdot)}$. Equation˜5 defines the marginal distribution of the clean data ($x_{0}=o$) which we denote as $p^{*}_{0}(x_{0})$ [^1] To adapt to the setting of diffusion, we use $x_{t}$ to denote the (masked) completions, and omit the prompt $q$.. To obtain the guidance at intermediate time steps $t>0$, we define the forward diffusion process for the target diffusion policy $\pi^{*}$ as following.
Definition 1.
The forward diffusion process of the target policy ($\pi^{*}$) satisfies $p^{*}_{t|0}(x_{t}|x_{0})=p_{t|0}(x_{t}|x_{0})$, where $p_{t|0}$ is the forward process of old diffusion policy $\pi_{\text{old}}$.
Since the reference diffusion policy is the initial policy, three policies have identical forward diffusion process, being $p^{*}_{t|0}(x_{t}|x_{0})=p_{t|0}(x_{t}|x_{0})=p^{\text{ref}}_{t|0}(x_{t}|x_{0})$, and thus, $p^{*}_{t|0}(x_{t}|x_{0})=p^{\prime}_{t|0}(x_{t}|x_{0})$, where $p^{\prime}$ is the geometric mixture diffusion $p^{\prime}_{t|0}(x_{t}|x_{0})\propto p_{t|0}(x_{t}|x_{0})^{\lambda}p^{\text{ref}}_{t|0}(x_{t}|x_{0})^{\beta}$. We can then obtain the energy guidance at all time step $t$.
Lemma 1 (Intermediate Energy Guidance on Discrete Diffusion).
The marginal probability distribution of the masked responses ($x_{t}$) in the diffusion process satisfies $p^{*}_{t}(x_{t})=p^{\prime}_{t}(x_{t})\cdot\exp\big(A_{t}(x_{t})\big)/Z_{t}$, which induces an energy-guided discrete diffusion:
$p^{*}_{0|t}(x_{0}|x_{t})\propto p^{\prime}_{0|t}(x_{0}|x_{t})\cdot\exp(A(x_{0})-A_{t}(x_{t})),$ (11)
where $A_{t}(x_{t})=-\log\mathbb{E}_{x_{0}\sim p^{\prime}_{0|t}(\cdot|x_{t})}[\exp\big(A(x_{0})\big)]$ is intermediate energy function for $t>0$, and $A(\cdot)$ is advantage function (Proof in ˜Lemma).
The guidance provided in Lemma˜1 demonstrates that it directs the sampling process toward generating completions that exhibit higher advantage values. However, conducting training-free guided sampling following Equation˜11 requires estimating the posterior mean of the exponential of advantage (Lu et al., 2023). Rather than relying on such estimation, we instead aim to find the training objective to directly approximate the target guided diffusion model.
Since existing masked dLLMs parametrize the concrete score (Meng et al., 2022), to apply the energy guidance, we aim to directly approximate target guided concrete score. Denote $x_{t}=(x_{t}^{1},\cdots,x_{t}^{d})$ and $\hat{x}_{t}$ is identical to $x_{t}$ except the $i$-th token is unmasked (i.e. $x_{t}^{i}=[M]$ and $\hat{x}_{t}^{i}\neq[M]$). Concrete score is defined as the marginal probability ratio between $\hat{x}_{t}$ and $x_{t}$:
$s(x_{t},t)\overset{\text{def}}{=}\frac{p(x_{t}^{1},\cdots,\hat{x}_{t}^{i},\cdots,x_{t}^{d})}{p(x_{t}^{1},\cdots,x_{t}^{i},\cdots,x_{t}^{d})}.$ (12)
We prove that the training objective to approximate the guided concrete score can be simplified as a weighted Denoising Concrete Score Matching (D-CSM) (Meng et al., 2022):
Theorem 1.
The model $s_{\theta}$ approximates the concrete score of the energy-guided discrete diffusion $p^{*}$ when the following loss objective is minimized. This objective is in a form of advantage-weighted Denoising Concrete Score Matching, which we call AW-D-CSM:
$\mathcal{L}_{\text{AW-D-CSM}}=$ $\mathbb{E}_{x_{0}\sim p^{\prime}_{0}(\cdot)}\bigg[\underbrace{\exp\big(A(x_{0})\big)}_{\text{Advantage Weight}}\cdot\underbrace{\mathbb{E}_{t\sim[0,T],p^{\prime}_{t|0}(x_{t}|x_{0})}[\|{s_{\theta}(x_{t},t)}-\frac{p^{\prime}_{0}(\hat{x}_{t}|x_{0})}{p^{\prime}_{0}(x_{t}|x_{0})}\|_{2}^{2}]}_{\mathcal{L}_{\text{D-CSM}}(x_{0})}\bigg].$ (13)
We provide the proof in Appendix Theorem. Additionally, D-CSM is an approximation of CSM (Meng et al., 2022), which is equivalent to Denoising score entropy (DSE) (Lou et al., 2024). For all $x_{0}$, it is satisfied up to multiplying a constant that $\mathcal{L}_{\text{D-CSM}}(x_{0})\Leftrightarrow\mathcal{L}_{\text{CSM}}(x_{0})\Leftrightarrow\mathcal{L}_{\text{DSE}}(x_{0})\Leftrightarrow\mathcal{L}_{\text{DCE}}(x_{0})$ (Ou et al., 2025). Therefore, AW-D-CSM can then be applied for both SEDD (Lou et al., 2024) and RADD (Ou et al., 2025) model such as LLaDA. Denote $p_{\theta}$ as the concrete score reparametrized model, AW-D-CSM can be converted to a weighted denoising cross-entropy loss (AW-DCE):
$\mathcal{L}_{\text{AW-DCE}}=$ $\mathbb{E}_{x_{0}\sim p^{\prime}_{0}(\cdot)}\bigg[\exp\big(A(x_{0})\big)\cdot\mathbb{E}_{t\sim[0,T],p^{\prime}_{t|0}(x_{t}|x_{0})}\big[\sum_{x^{i}_{t}=[\texttt{mask}]}-\frac{1}{t}\log p_{\theta}(x_{0}^{i}|x^{\text{UM}}_{t})\big]\bigg].$ (14)
DSE and DCE objectives both can be used for likelihood approximation in fine-tuning (Ou et al., 2025; Nie et al., 2025a; Yang et al., 2025) since they can serve as negative ELBO (Lou et al., 2024; Shi et al., 2025). Thus, the advantage-weighted objective AW-DCE (or AW-DSE) used to learn energy-guided score is in a weighted log-likelihood form as in wd1 with only $w^{+}$ (i.e. WLL loss in Equation˜6), which contributes to our main theoretical findings:
Remark 1.
In the context of applying RL to masked discrete diffusion, the advantage-weighted log-likelihood (WLL) objective (Equation˜6) induced by reverse-KL policy optimization, is equivalent to the objective of training energy-guided diffusion models, where the energy function is the negative advantage. Formally, $\mathcal{L}_{\text{WLL}}\Leftrightarrow\mathcal{L}_{\text{AW-DCE}}$ when DCE is used for likelihood approximation.
Remark 2.
Additionally, based on DCE likelihood, the additional penalty term on negative samples used to extend WLL to wd1 loss can be viewed as applying data unlearning by minimizing the ELBO (Alberti et al., 2025), where the data $\{x^{-}_{0}\}$ (negative samples) has probability distribution $p_{\text{data}}(x^{-}_{0})\propto p^{\prime}_{0}(x^{-}_{0})\exp(-A(q,x^{-}_{0}))$, which corresponds to a Boltzmann distribution that places higher probability mass on regions with lower advantage values (more details in Section˜D.1).
@@ 5 Experiments
    Improved reasoning capabilities than existing methods on popular reasoning benchmarks;
@@
    reduced computational burden, as reflected by decreased runtime, lower FLOPs and numbers of function evaluations (NFEs) per training step, number of training steps and rollouts; and
@@
Experimental Setup. We perform reinforcement learning (RL) fine-tuning on the LLaDA-8B-Instruct model (Nie et al., 2025a) with Low-Rank Adaptation (LoRA) on: GSM8k (Cobbe et al., 2021), MATH (Lightman et al., 2023), Sudoku (Arel, 2025), and Countdown (Pan et al., 2025). As for decoding, we follow the default strategy Mounier and Idehpour (2025); Arriola et al. (2025); Wang et al. (2025b). Our main baseline is d1 (Zhao et al., 2025), the first RL method developed for masked diffusion LLMs (dLLMs). We reproduce the baseline methods Diffu-GRPO, which applies diffusion-based GRPO training directly to the LLaDA base model, and d1, which performs SFT before applying Diffu-GRPO. We use s1K  (Muennighoff et al., 2025) data for SFT in d1. We also compare with SDPO (Han et al., 2025), TCR (Wang et al., 2025d), and MDPO (He et al., 2025) on benchmarks GSM8K and MATH500. MDPO is reproduced based on the official implementation and the training dataset (He et al., 2025).
Implementation. As for wd1, we conduct training on the same dataset as in d1 (Zhao et al., 2025): training splits on GSM8k and MATH, and the dataset splits provided by Zhao et al. (2025) on Sudoku and Countdown. In our implementation of wd1, we apply the same likelihood approximation method as d1. The hyperparameters used in our method and our reproduction of d1 are listed in Table˜6 and Table˜5. As for wd1++, we train on dataset provided by (He et al., 2025), which is sampled from OpenR1 dataset (Face, 2025). Since previous works (Yu et al., 2025) have demonstrated that the reference policy is empirically unnecessary, we set $\beta=0$ and $\lambda=1$ to eliminate $\pi_{\text{ref}}$ in practice. We report results using zero-shot and pass@1 evaluation on sequence lengths of 256 and 512 tokens.
Table 1: Test Accuracy (%) of wd1 and d1. We reproduce d1 and vary completion length. Our approach without SFT, demonstrates particularly higher accuracy on Sudoku [^3] In the technical report version of this work, our method achieved scores of 25.2 and 24.2 on Sudoku after 5K training steps. In this paper, we extend the training to 12.5K steps, and wd1 results in improved performance.and Countdown.
Model / Gen Len Sudoku Countdown GSM8K MATH500
@@
+ wd1 76.4 62.8 51.2 46.1 80.8 82.3 34.4 39.0
[…]
Method SFT RL Training
@@
wd1 0 hrs 81.16 sec/step $8.887\times 10^{15}$/step $\mu$/step
@@ 5.1 Main Results
Superior Reasoning Ability. In Table˜1, we observe that wd1, even without supervised fine-tuning or using any supervised data, consistently outperforms our reproduced implementation of d1. Notably, wd1 surpasses d1 by 43% in test accuracy on the Sudoku task, and achieves up to a 25% improvement on Countdown with maximum length 256. Relative to the base LLaDA model, the performance gain reaches as high as 54% on Sudoku and 42% on Countdown. On math problem-solving benchmarks GSM8K and MATH500, wd1 attains slightly higher accuracy. Nevertheless, the extended method wd1++ obtains significantly better accuracy. In Table˜3 (left), we further compare with concurrent baselines released in recent months. wd1++ outperforms the baselines including strong one MDPO.
Reduced Training Cost. Table˜2 demonstrates that the training cost required by wd1 is substantially lower than that of d1. Unlike d1, wd1 does not require a SFT stage, which alone accounts for approximately two hours of training in d1. wd1 achieves additional speedup during the RL phase, where runtime is measured by averaging over $\mu=8$ inner gradient steps per global step. Notably, the time efficiency gap is expected to widen further under settings with larger maximum sequence lengths and more diffusion steps. This efficiency gain is further supported by a reduced FLOPs and number of function evaluations (NFEs) per step, as wd1 bypasses the need to approximate the likelihood of the old policy. We exclude NFEs associated with sampling, since both methods share identical sampling costs as wd1 removes the reference policy regularization.
In Table˜3 (right), we report the training cost required to obtain the best post-trained models on GSM8K and MATH500, measured in terms of the number of training steps and rollouts. wd1++ requires $10\times$ fewer rollouts to achieve superior performance, clearly demonstrating the efficiency of our method. This rapid convergence arises primarily from the exponential advantage weights applied to the log-likelihood in wd1. In contrast, standard RL methods such as GRPO and PPO weight the log-likelihood (or policy ratio) terms directly by the advantage function.
Table 3: Left: Extended method wd1++ compared to concurrent RL methods to fine-tune LLaDA-8B-Instruct. Methods denoted by “(full)” perform full fine-tuning. Right: Training cost to obtain the best model on GSM8K and MATH500. We count the total number of steps of policy iteration (model weights update), and the number of rollouts used for training (see Table 8 for details on counting).
Model GSM8K MATH500
LLaDA-8B-Instruct 78.2 36.2
+ diffu-GRPO (Zhao et al., 2025) 80.7 39.0
+ d1 (Zhao et al., 2025) 82.0 38.0
+ SDPO (Han et al., 2025) (full) 81.2 -
+ TCR (Wang et al., 2025d) 83.0 41.4
+ MDPO (He et al., 2025) (full) 83.4 43.4
+ wd1 82.3 39.0
+ wd1 (full) 82.7 43.6
+ wd1++ (full) 84.5 44.2
@@ 5.2 Ablation Study
Table 4: Ablation on SFT and Negative Samples Weight ($w^{-}$). We conduct wd1 training after SFT (wd1-SFT) and with only $w^{+}$ (namely wd1-P or WLL defined in Equation˜6) [^5] To facilitate efficient ablation studies, we restrict our comparisons to checkpoints saved prior to 5000 steps.. Results show that wd1 performs better without SFT on planning and math tasks. Removing negative sample reinforcement ($w^{-}$) significantly hurts performance, highlighting its importance.
Model / Gen Len Sudoku Countdown GSM8K MATH500
@@
We present an ablation study in Figure 4. Notably, we observe that supervised fine-tuning (SFT) yields only marginal improvements within our approach, with a slight gain in the Sudoku task. This contrasts with d1, where SFT plays a significant role in improving performance. These findings indicate that wd1 can eliminate the need for an SFT phase, thereby simplifying the training pipeline and substantially reducing computational cost. Additionally, we evaluate the impact of removing the negative-weighted term by setting $w^{-}=0$, thus relying solely on the positive advantage weights $w^{+}$. We provide further ablation on the combined method between $w^{+}$ and $w^{-}$ in Table˜9. The results highlight the importance of explicitly penalizing the likelihood of low-advantage completions, thereby reinforcing the role of negative samples, and emphasize the critical balance between the two weights.
Figure 2: Training rewards of wd1 under different combined weights on Sudoku.
We further assess sensitivity to the relative weighting of positive and negative samples. The combined weight (cw) corresponds to $\lambda$ in the mixture $\lambda w^{+}+(1-\lambda)w^{-}$, which scales the log-likelihood term in wd1. Training on negative samples alone (cw=0.0=0.0) yields a pronounced deterioration in performance relative to our default setting (cw=0.5=0.5). The results reinforce our argument that a balanced contribution of positive and negative weights is most effective. In the absence of positive samples, the reinforcement-learning signal collapses and optimisation becomes largely ineffective. A large emphasis on positive samples (cw=0.8=0.8) causes performance to deteriorate more rapidly, highlighting the critical role of negative samples in weighted log-likelihood methods.